# Self-supervised Learning for Geospatial AI: A Survey

### v1

@@ Abstract
The proliferation of geospatial data in urban and territorial environments has significantly facilitated the development of geospatial artificial intelligence (GeoAI) across various urban applications. Given the vast yet inherently sparse labeled nature of geospatial data, there is a critical need for techniques that can effectively leverage such data without heavy reliance on labeled datasets. This requirement aligns with the principles of self-supervised learning (SSL), which has attracted increasing attention for its adoption in geospatial data. This paper conducts a comprehensive and up-to-date survey of SSL techniques applied to or developed for three primary data (geometric) types prevalent in geospatial vector data: points, polylines, and polygons. We systematically categorize various SSL techniques into predictive and contrastive methods, discussing their application with respect to each data type in enhancing generalization across various downstream tasks. Furthermore, we review the emerging trends of SSL for GeoAI, and several task-specific SSL techniques. Finally, we discuss several key challenges in the current research and outline promising directions for future investigation. By presenting a structured analysis of relevant studies, this paper aims to inspire continued advancements in the integration of SSL with GeoAI, encouraging innovative methods to harnessing the power of geospatial data.
Index Terms:
@@
Geospatial artificial intelligence, self-supervised learning, spatial representation learning, contrastive learning.
I Introduction
@@
The digitization of urban and territorial environments has significantly enhanced the collection of extensive data in geospatial space, including user check-ins, traffic records, building footprints, trajectories, etc. Such data serves as the foundation of smart city applications, such as spatial keyword search [1, 2], location-based services [3, 4], geospatial knowledge graph [5], intelligent transportation systems [6, 7], and socioeconomic indicator prediction [8, 9]. This vast pool of data, while rich in potential, has long faced limitations and challenges for analysis and utilization for the development of geospatial artificial intelligence (GeoAI) [10], especially for the advanced models in the context of deep learning. In particular, the models are usually trained on specific tasks with abundant labeled data (e.g., traffic records), which can be both limited and costly to be accessible, especially considering data privacy regulations [11]. Additionally, despite the commonalities shared by various urban tasks, such as the close relationship between population density and land use, these models usually suffer from limited generalization across downstream tasks due to their reliance on task-specific supervision signals.
@@
In recent years, self-supervised learning (SSL) [12] has emerged as a paradigm that can reduce the dependency on annotated labels while producing generic and task-agnostic data representations. The core principle of SSL is to extract transferrable knowledge from the target data through well-designed self-supervised tasks (i.e., pretext tasks), wherein the supervision signals are automatically generated from the data itself. SSL has achieved notable success across various domains and diverse data modalities, including images [13, 14], videos [15], language [16, 17], graphs [18, 19], time series [20], etc. For example, massive text corpora are structured in an autoregressive way, which is suitable for next token prediction for the training of large language models (LLMs) [21]. Images can be processed through data augmentation operations to produce multiple views, with the models trained to maintain invariance across views [13].
@@
A primary motivation of applying SSL techniques for GeoAI is to learn effective and generalizable representations (embeddings) for various forms of geospatial objects, such as POI, road segments, and urban regions. These objects underpin a variety of human activities within urban environments, and therefore serve as fundamental analytical units for various urban analytical tasks. For example, road networks support most human mobility activities within a city. As a result, a range of tasks, such as the predictions of traffic speed, congestion, and destination, commonly regard individual road segments as analytical units. In this case, SSL serves as a compelling framework to derive general-purpose representations for each type of geospatial data, capturing both the properties of each object as well as the complex interplay among different objects. The derived representations can be readily utilized to train simpler models for various downstream tasks while maintaining effective performance.
Apart from the strong generalization capabilities, the interest of applying SSL techniques is also driven by their intrinsic advantages to operate without extensive labeled datasets. This characteristic aligns seamlessly with addressing a common challenge in the geospatial domain—the scarcity of labeled data. Consequently, SSL offers a viable alternative to the conventional approach of developing specific deep learning models for each urban analytical task with sufficient labeled data in a supervised learning paradigm. However, geospatial data, embedded within geographical space, presents various forms with unique spatial features and adheres to certain geographical principles, such as the Laws of Geography [22]. As a result, SSL techniques commonly applied in other domains often fall short in capturing the spatial semantics attached to geospatial data. Furthermore, certain SSL operations utilized in other domains, such as data augmentation, need to be carefully designed to reflect the characteristics of different geospatial objects.
@@
Despite the clear distinction and a growing body of literature, the application of SSL within the domain of GeoAI, remains insufficiently discussed and explored. To bridge this gap, this survey provides a comprehensive and systematic review of the up-to-date SSL techniques tailored for geospatial data. In particular, we focus on three primary data types depending on their geometric forms, i.e., points, polylines, and polygons, which are prevalent in urban contexts. We organize a structured way to present the specialized SSL studies related to these data types, focusing particularly on those that operate independently of specific primary tasks and supervised settings. Furthermore, considering the rapid development in this domain, we discuss several emerging trends that have recently appeared, including multi-type learning and adaptation to language space. We also provide a brief review of task-specific SSL techniques applied to geospatial data. This survey aims to cover a wide range of model scopes, from specialized SSL models to advanced multi-modality models, applied to different geospatial data types. The main contributions of this survey are summarized as follows:
@@
  • We present a detailed and up-to-date review of SSL techniques for geospatial data, focusing on three types of geospatial vector data in urban environments: points, polylines, and polygons. To the best of our knowledge, this work is the first to systematically discuss SSL techniques tailored for geospatial data.
@@
  • We introduce a comprehensive and structured taxonomy for specialized SSL models designed for the studied geospatial data types. Our categorization includes an analysis of intrinsic characteristics and contextual information within each data type, encompassing discussion on both predictive and contrastive SSL implementations.
@@
  • We review several recent advancements and task-specific SSL techniques for geospatial data, providing insights into the emerging and promising trends applied to GeoAI.
@@
  • We discuss several key challenges for SSL in GeoAI, and propose promising future directions to advance this domain.
@@
Related Surveys and Our Distinction: Several recent surveys have discussed the application of SSL, mainly focusing on other domains, such as general SSL [12], SSL for computer vision [14], graph data [18, 19], time series [20], and recommender systems [23]. Recognizing the lack of dedicated SSL exploration within the domain of GeoAI, this paper presents a comprehensive and systematic review of SSL tailored for geospatial data, which serves as the foundations in numerous downstream applications. On the other hand, several recent surveys have paid attention to spatio-temporal data analytics with various emphases, such as trajectory data mining [24, 25], urban foundation models [26], geo-location encoding [27], and supervised/generative deep learning [28, 29]. These surveys generally fall within the domain of GeoAI, and may overlap with some studies discussed in this paper. However, this survey is distinctive as it pioneers in structuring the knowledge from the unique perspective of SSL, detailing its developments across different geospatial data types and providing novel insights into the applications of SSL in geospatial contexts.
@@
Paper Structure: The rest of this survey is organized as follows. Section II provides definitions, preliminary concepts, and background knowledge necessary for the subsequent sections. Section III, IV, and V look into the details of specialized SSL techniques applied to geospatial data with three distinct geometric types: points, polylines and polygons, respectively. For each data type, representative data instances are further elaborated in the context of SSL, including POI for points, trajectory and road network for polylines, and region for polygons. Section VI highlights emerging trends in this domain, such multi-type learning and adaptation to language space, and provides an overview for task-specific SSL techniques for geospatial data. Section VII discusses several promising future research directions. Finally, Section VIII concludes this paper.
II Definition and Background
@@
In this section, we first introduce the definition of three types of geospatial vector data examined in this paper. Then we present the paradigms of SSL applied to geospatial objects based on the traits of pretext tasks. Last, we discuss some preliminaries on typical models that encode these geospatial data types.
II-A Definition of Geospatial Data Types
The diverse and rich data sources collected in urban spaces, which display unique geospatial characteristics, can be mainly categorized into three data types in terms of their geometric representations: points, polylines and polygons [30]. These data types are illustrated in Fig. 1. Then we provide formal definitions for each of these types to establish a clear understanding necessary for subsequent discussions and analyses.
@@ II-A Definition of Geospatial Data Types
A point in the geospatial context is represented as $p=(l,x)$, where $l$ denotes the geographical coordinates, and $x$ refers to the associated features of the point, such as attributes or readings. This data type indicates the spatial locations equipped with contextual information, applicable to data instances such as POIs and sensor measurements.
@@
A polyline in the geospatial context is defined as a sequence of connected line segments, represented by a list of vertices $\mathcal{L}=[(l_{1},x_{1}),...,(l_{n},x_{n})]$, where $l_{i}$ denotes the geographical coordinates of the $i$-th point, and $x_{i}$ denotes its associated features, such as timestamps or semantic tags. This data type captures the sequential nature and directionality of spatial paths, applicable to data instances such as trajectories and road networks.
@@
A polygon in the geospatial context is defined as a closed shape consisting of a sequence of line segments that connect to enclose an area, denoted by $\mathcal{B}=[(l_{1},l_{2}),...,(l_{n-1},l_{n}),(l_{n},l_{1})]$, where $l_{i}$ represents the geographical coordinates of the $i$-th vertex. The vertices are sequentially linked, with the last vertex reconnecting to the first vertex to complete the polygon. By default, the line segments of a polygon are arranged to ensure that it does not intersect with itself, maintaining a simple, closed loop. In urban contexts, this data type can be employed to describe administrative regions or subzones in urban spaces. Therefore, research on polygons often extends its focus beyond its geometric configuration, placing emphasis on the semantic patterns exhibited by objects within its enclosed area.
II-B Paradigms for Self-supervised Learning
SSL techniques for geospatial objects can be divided into two categories: predicative and contrastive, according to the design of pretext tasks.
II-B1 Predicative Methods
Predictive methods employ pretext tasks that are formulated as prediction problems, with objectives derived from the original data instances. Specifically, these methods involve tasks like the reconstruction of corrupted geospatial objects using a subset of available data, or the prediction of auxiliary labels that are extracted from the attributes or structures of geospatial objects. They can be formulated as:
@@ II-B1 Predicative Methods
where $f_{\theta}$ and $p_{\phi}$ represent the geospatial encoder and the pretext decoder respectively. $\mathcal{D}$ denotes the target geospatial objects, which can be of any data types. $\mathcal{D}_{c}$ is the context information associated with $\mathcal{D}$, and $\mathcal{D}_{t}$ denotes the prediction objectives, which could either be the original data instance for reconstruction tasks, or additional features excluded from the encoder input for auxiliary label prediction. $\mathcal{L}_{pre}$ is the loss function that measures the prediction error, such as cross-entropy loss or mean squared error (MSE) loss.
II-B2 Contrastive Methods
@@
where $f_{\theta}$ and $p_{\phi}$ represent the geospatial encoder and the pretext decoder respectively. $\mathcal{\tilde{D}}^{1}$ and $\mathcal{\tilde{D}}^{2}$ are two distinct views generated from the target geospatial objects $\mathcal{D}$, which can belong to any geospatial data types, and $\mathcal{\tilde{D}}_{c}^{1}$ and $\mathcal{\tilde{D}}_{c}^{2}$ are the context information associated with these respective views. $\mathcal{L}_{con}$ is the contrastive loss function that quantifies the degree of agreement, typically measured by mutual information estimator [12], such as InfoNCE [31], JS divergence [32] and triplet loss [33].
II-C Preliminaries on Geospatial Encoder
Given the diversity of geospatial data types discussed in this survey, each with its unique geometric (locational) and intrinsic properties, the utilized geospatial encoder $f_{\theta}$ would be varied to accommodate their distinct characteristics. Therefore, we provide a brief introduction on several neural network modules that are frequently employed or adapted as geospatial encoders.
II-C1 Graph Neural Networks

### v2

# Self-Supervised Representation Learning for Geospatial Objects: A Survey

@@ Abstract
The proliferation of various data sources in urban and territorial environments has significantly facilitated the development of geospatial artificial intelligence (GeoAI) across a wide range of geospatial applications. However, geospatial data, which is inherently linked to geospatial objects, often exhibits data heterogeneity that necessitates specialized fusion and representation strategies while simultaneously being inherently sparse in labels for downstream tasks. Consequently, there is a growing demand for techniques that can effectively leverage geospatial data without heavy reliance on task-specific labels and model designs. This need aligns with the principles of self-supervised learning (SSL), which has garnered increasing attention for its ability to learn effective and generalizable representations directly from data without extensive labeled supervision. This paper presents a comprehensive and up-to-date survey of SSL techniques specifically applied to or developed for geospatial objects in three primary vector geometric types: Point, Polyline, and Polygon. We systematically categorize various SSL techniques into predictive and contrastive methods, and analyze their adaptation to different data types for representation learning across various downstream tasks. Furthermore, we examine the emerging trends in SSL for geospatial objects, particularly the gradual advancements towards geospatial foundation models. Finally, we discuss key challenges in current research and outline promising directions for future investigation. By offering a structured analysis of existing studies, this paper aims to inspire continued progress in integrating SSL with geospatial objects, and the development of geospatial foundation models in a longer term.
keywords:
@@
Geospatial artificial intelligence, Spatial data mining , Self-Supervised learning , Spatial representation learning , Geospatial foundation models
journal: Information Fusion
[…]
[a]organization=College of Computing and Data Science, Nanyang Technological University,country=Singapore
[…]
[b]organization=Department of Physical Geography and Ecosystem Science, Lund University,country=Sweden
[…]
[c]organization=School of Computer Science, University of Auckland,country=New Zealand
1 Introduction
@@
The digitization of urban and territorial environments has significantly expanded the collection of extensive geospatial data associated with various objects on our planet, including road segments, buildings, neighborhoods, etc. The vast repository of data serves as the foundation of smart city applications, such as spatial keyword search [1, 2], location-based services [3, 4], geospatial knowledge graph [5], intelligent transportation systems [6, 7], and socioeconomic indicator prediction [8, 9]. Despite the richness of geospatial data, its effective mining and utilization remain challenging. In particular, a fundamental limitation arises from the task-specific nature of many deep learning models. These models are typically trained using supervised learning in specific tasks with abundant domain-specific labeled datasets (e.g., traffic records), which can be limited or costly to acquire and restricted by data privacy regulations [10]. Additionally, while various urban tasks exhibit intrinsic commonalities, such as the close relationship between population density and land use patterns, these models usually suffer from limited generalization across downstream applications due to their reliance on task-specific supervision signals and specialized model architectures. This lack of adaptability restricts the broader applicability of previous approaches across different geospatial analytics tasks, which entails the need for more flexible and generalizable learning paradigms for a variety of geospatial analyses.
@@
In response to these challenges, self-supervised learning (SSL)  [11] has emerged in recent years as a promising paradigm that reduces the dependency on annotated labels while producing task-agnostic and general-purpose data representations. The core principle of SSL is to extract transferrable knowledge from the target data through well-designed self-supervised tasks (i.e., pretext tasks), wherein the supervision signals are automatically generated from the data itself. SSL has achieved notable success across various domains and diverse data modalities, including images [12, 13], videos [14], language [15, 16], graphs [17, 18], time series [19], etc. For example, massive text corpora are structured in an autoregressive generation framework, which is well-suited for next token prediction for the training of large language models (LLMs) [20]. In addition, images are processed through data augmentation operations to produce multiple views, with models trained to produce invariant representations across views for computer vision tasks [12].
@@
A key motivation for applying SSL techniques in the geospatial domain is to learn effective and generalizable representations (embeddings) for various forms of geospatial objects, such as points-of-interest (POI), road segments, and urban regions. These objects underpin a variety of human activities within urban environments, and therefore serve as fundamental analytical units for numerous urban analytical applications. For example, road networks are critical infrastructures that support human movement activities within a city. As a result, a variety of tasks, such as the prediction of traffic speed, congestion, and destination, commonly regard individual road segments as analytical units. In these scenarios, SSL offers a powerful framework for deriving general-purpose representations of geospatial objects through fusing information from multiple perspectives, capturing both the intrinsic characteristics of each object and the complex interplay among different objects. The representations learned through SSL can be readily leveraged to train simpler models (e.g., a linear model) for various downstream tasks while maintaining effective performance.
Apart from the strong generalization capabilities, the interest in applying SSL techniques is also driven by the ability to work without extensive labeled datasets. This advantage addresses a long-standing challenge in the geospatial domain– the scarcity of task-specific labeled data. As a result, SSL presents a viable alternative to the conventional supervised learning paradigm, which necessitates the development of specialized deep learning models trained on sufficiently labeled data for each downstream application. By leveraging the self-supervised signals derived from data itself, SSL has the potential to better tackle diverse geospatial (urban) analytical applications.
However, geospatial objects, embedded within geographic spaces, exhibit various forms and spatial attributes while adhering to geographical principles, such as the Laws of Geography [21]. These spatial characteristics introduce significant challenges when applying standard SSL techniques developed for other domains, as they often fail to capture the intricate spatial semantics attached to geospatial objects. Besides, certain SSL components, such as data augmentation and pretext tasks, must be carefully designed and adapted to preserve the spatial and structural integrity of geospatial objects. Unlike in vision or language domains, where augmentations such as cropping, rotation, or synonym replacement are widely used, augmentations for spatial objects must ensure consistency with geographical relationships and dependencies. Moreover, geospatial objects are often associated with heterogeneous context information, capturing diverse yet complementary aspects of their intrinsic properties. Integrating this information requires the development of effective modeling and fusion strategies to produce high-quality data representations. Considering these unique challenges in the geospatial domain, recent research has introduced novel SSL techniques that incorporate spatial awareness and domain-specific constraints, ensuring the effective adaptation of SSL within the geospatial context.
@@
Despite the growing body of literature, the application of SSL within the geospatial domain remains insufficiently discussed and summarized. To bridge this gap, this survey provides a comprehensive and systematic review of up-to-date SSL techniques tailored or developed for learning geospatial object representations, which in turn facilitate various geospatial analyses. In particular, we focus on three primary geospatial data types, categorized based on their geometric forms: points, polylines, and polygons. We adopt a structured framework to present the specialized SSL studies for these data types, focusing particularly on methods that operate independently of specific tasks and supervised settings. For each data type, we analyze how SSL techniques encode intrinsic attributes while integrating heterogeneous context information associated with geospatial objects. Besides, we review studies that utilize SSL to fuse multiple geospatial data types for learning representations. Building on the insights from the surveyed studies, we identify and analyze emerging trends in the development of geospatial foundation models. Furthermore, since SSL techniques serve as a universal paradigm that can also be utilized to enhance the supervised models as auxiliary objects in geospatial applications such as spatio-temporal forecasting, we provide a brief overview of task-specific SSL implementations applied to several domain-specific scenarios, supplementing the research landscape of SSL in this field. This survey aims to cover a wide range of model scopes, from specialized SSL models to advanced foundation models, applied to different geospatial data types and analyses. The main contributions of this survey are summarized as follows:
@@
  • We present a detailed and up-to-date review of SSL techniques for geospatial objects, focusing on three types of geospatial data, mainly in urban environments: Point, Polyline, and Polygon (Region). To the best of our knowledge, this work is the first to systematically discuss SSL techniques tailored for learning representations in the geospatial domain.
@@
  • We introduce a comprehensive and structured way for specialized SSL models designed for the studied data types. Our categorization includes an analysis of intrinsic characteristics and heterogeneous context information within each data type, encompassing discussion on both predictive and contrastive SSL implementations.
@@
  • We review several recent advancements based on foundation models and task-specific SSL techniques for geospatial data, providing insights into the emerging trends.
@@
  • We discuss several key challenges for SSL in geospatial domain, and propose potential future directions to advance the related research.
@@
Related Surveys and Our Distinction: Several recent surveys have discussed the application of SSL, mainly focusing on other domains, such as general SSL [11], SSL for computer vision [13], graphs [17, 18], time series [19], and recommender systems [22]. However, despite SSL’s growing importance in data-driven analytics, its adaptation in the geospatial domain remains underexplored. Recognizing this limitation, our survey presents a comprehensive and systematic review of SSL tailored for geospatial data, which serves as the foundations in numerous downstream applications to enhance urban intelligence. On the other hand, several recent surveys have paid attention to spatio-temporal data analytics with varying emphases, such as trajectory data mining [23, 24], urban foundation models [25], geo-location encoding [26], and supervised or generative deep learning [27, 28]. While these works contribute to the broader field of GeoAI and may overlap with some studies discussed in this paper, our survey distinguishes itself by structuring knowledge from the unique perspective of SSL. Specifically, we provide a detailed summary of SSL developments across different geospatial data types, systematically categorizing existing methodologies and synthesizing their implementation in geospatial contexts.
@@
Paper Structure: The rest of this survey is organized as follows. Section 2 provides definitions, preliminary concepts, and background knowledge necessary for the subsequent sections. Section 34, and 5 look into the details of specialized SSL techniques applied to data with three distinct geometric types: points, polylines and polygons, respectively. For each data type, representative data instances are further elaborated in the context of SSL, including POI for points, trajectory and road network for polylines, and region for polygons. Section 6 presents SSL techniques for the integration of multiple data types. Section 7 presents our discussion on the emerging trends based on the development of geospatial foundation models, and provides an overview for task-specific SSL techniques for geospatial data. Section 8 provides several promising future research directions. Finally, Section 9 concludes this paper.
2 Definition and Background
@@
In this section, we introduce the definition of three types of geospatial vector data examined in this paper. Then we present the paradigms of SSL applied to geospatial objects based on the traits of pretext tasks. Last, we discuss preliminaries on typical models that encode these geospatial data types.
2.1 Definition of Geospatial Data Types
In this survey, we adopt the widely recognized classification scheme in spatial database research, which categorizes geospatial objects by their geometric representations into three types: Point, Polyline, and Polygon (Region) [29]. This scheme is well-suited to our study, as each type corresponds to distinct real-world geospatial phenomena that require different representation strategies. The three data types are illustrated in Figure 1. The formal definitions of the three data types are provided as follows.
@@ 2.1 Definition of Geospatial Data Types
A geospatial point object is represented as $p=(l,x)$, where $l$ denotes the geographical coordinates, and $x$ refers to the associated features of the point, such as attributes or readings. This data type indicates the spatial locations equipped with contextual information, applicable to data instances such as POIs and sensor measurements.
@@
A geospatial polyline object is defined as a sequence of connected line segments, represented by a list of vertices $\mathcal{L}=[(l_{1},x_{1}),...,(l_{n},x_{n})]$, where $l_{i}$ denotes the geographical coordinates of the $i$-th point, and $x_{i}$ denotes its associated features, such as timestamps or semantic tags. This data type captures the sequential nature and directionality of spatial paths, applicable to data instances such as trajectories and road networks.
@@
A geospatial polygon object is defined as a closed shape (region) consisting of a sequence of line segments that connect to enclose an area, denoted by $\mathcal{B}=[(l_{1},l_{2}),...,(l_{n-1},l_{n}),(l_{n},l_{1})]$, where $l_{i}$ represents the geographical coordinates of the $i$-th vertex. The vertices are connected sequentially, with the first vertex repeated as the last to close the polygon. By default, the line segments of a polygon are arranged to ensure that it does not intersect with itself, maintaining a simple, closed loop. This data type can be employed to describe administrative regions or subzones in urban spaces. Therefore, research on polygons often extends its focus beyond its geometric configuration, placing emphasis on the semantic patterns exhibited by objects within or close to its enclosed area.
2.2 Paradigms for Self-supervised Learning
Geospatial data instances in practical scenarios can be conceptualized as a combination of geometric forms with intrinsic attributes (e.g., spatial coordinates) and associated context information (e.g., textual content). Therefore, SSL process in geospatial context aims to integrate the geometric form of the target geospatial object and its contextual information to derive effective representations. These representations are learned through the training objectives (pretext tasks), where supervision signals are automatically extracted from the geospatial data itself, eliminating the need for additional label annotations. Based on the design of pretext tasks, SSL techniques for geospatial objects can be divided into two categories: predictive and contrastive methods.
2.2.1 Predictive Methods
Predictive methods employ pretext tasks that are formulated as prediction problems, with objectives derived from the original data instances. Specifically, these methods involve tasks like the reconstruction of corrupted geospatial objects using a subset of available data, or the prediction of auxiliary labels that are readily extracted from the attributes or structures of geospatial objects. They can be formulated as:
@@ 2.2.1 Predictive Methods
where $f_{\theta}$ and $p_{\phi}$ represent the geospatial encoder and the pretext decoder respectively. The geospatial encoder, which is introduced in Section 2.3, is responsible for deriving representations for geospatial objects, while the pretext decoder is typically a lightweight structure, such as a shallow multi-layer perceptron (MLP), to map the representations to the space of the prediction objectives. $\mathcal{D}$ denotes the target geospatial objects, which can be of any data types. $\mathcal{D}_{c}$ is the context information associated with $\mathcal{D}$, and $\mathcal{D}_{t}$ denotes the prediction objectives, which could either be the original data instance for reconstruction tasks, or additional features excluded from the encoder input for auxiliary label prediction. $\mathcal{L}_{pre}$ is the loss function that measures the prediction error, such as cross-entropy loss or mean squared error (MSE) loss.
2.2.2 Contrastive Methods
@@
where $f_{\theta}$ and $p_{\phi}$ represent the geospatial encoder and the pretext decoder respectively. The pretext decoder is usually a projection head for linear transformation [12]. $\mathcal{\tilde{D}}^{1}$ and $\mathcal{\tilde{D}}^{2}$ are two distinct views generated from the target geospatial objects $\mathcal{D}$, which can belong to any geospatial data types, and $\mathcal{\tilde{D}}_{c}^{1}$ and $\mathcal{\tilde{D}}_{c}^{2}$ are the context information associated with these respective views. $\mathcal{L}_{con}$ is the contrastive loss function that quantifies the degree of agreement, typically measured by mutual information estimator [11], such as InfoNCE [30], JS divergence [31] and triplet loss [32].
2.3 Preliminaries on Geospatial Encoder
Given the diversity of geospatial data types discussed in this survey, each with its unique geometric (locational) and intrinsic properties, the utilized geospatial encoder $f_{\theta}$ would be varied to accommodate their distinct characteristics. Therefore, we provide a brief introduction on several neural network modules frequently employed or adapted as geospatial encoders.
2.3.1 Graph Neural Networks